Find the Reynolds number if a fluid of viscosity 0.4 Ns/m2 and relative density of 900 kg/m3
through a 20 mm pipe with a velocity of 2.5 m/s. Is the flow Laminar or Turbulent?
Note: If Reynolds < 2000 the flow is laminar
If Reynolds > 4000 the flow is turbulent
If 2000 < Re < 4000 it is transitional flow
Re =
1
Expert's answer
2020-06-02T17:51:47-0400

We can calculate the Reynolds number as (see https://en.wikipedia.org/wiki/Reynolds_number)

"\mathrm{Re} = \dfrac{\rho v D}{\mu} = \dfrac{900\,\mathrm{kg/m^3}\cdot2.5\,\mathrm{m/s}\cdot0.020\,\mathrm{m}}{0.4\,\mathrm{Ns/m^2}} = 112.5."

112.5 < 2000, therefore the flow is laminar.